ABC Inc. had current earnings and profits of $50,000 when it distributed to an individual shareholder land that the corporation held as an investment. On the date the land was distributed, ABC Inc.’s adjusted basis in the land was $10,000, the fair market value of the land was $50,000, and the land was encumbered by a $30,000 mortgage, which liability was assumed by the shareholder. There were no other transactions that might affect ABC Inc.’s earnings and profits for the year. What was the amount of ABC Inc.’s earning and profits at the end of the year?

Group of answer choices

$30,000

$50,000

$60,000

$70,000
Answer #1

Given data

ABC Inc. had current earnings and profits of $50,000 when it distributed to an individual shareholder land that the corporation held as an investment. On the date the land was distributed, ABC Inc.’s adjusted basis in the land was $10,000, the fair market value of the land was $50,000, and the land was encumbered by a $30,000 mortgage, which liability was assumed by the shareholder.

.

Before adjustment ABC Inc. Has earnings and profits of $50,000

.

Net decrease from current earnings and profits would be different between fair market value of the land and amount of mortgage debt

fair market value of the land = 50000

  amount of mortgage debt = 30000

Difference = 50000 - 30000 = 20000

.

This difference shall be treated as dividend paid ans should deduct from current earnings and profits of $50000

.

What was the amount of ABC Inc.’s earning and profits at the end of the year?

.

current earnings and profits of $50,000

Less: Net decrease = $20000

earning and profits at the end of the year = 50000 - 20000 = $30000

.

Ans: $30,000